One drain can empty a storage tank in 8.5 hours. When this drain is used along with a second drain, the tank can be emptied in 5 hours. How long(to the nearest 0.1 hour) would it take the second drain, working alone, to empty the tank?
Answer by stanbon(75887)

One drain can empty a storage tank in 8.5 hours. When this drain is used along with a second drain, the tank can be emptied in 5 hours. How long(to the nearest 0.1 hour) would it take the second drain, working alone, to empty the tank?
--------
1st drain DATA:
time = 8.5 hrs/job ; rate = 1/8.5 job/hr
------------------
together drain DATA:
time = 5 hrs/job ; rate = 1/5 job/hr
-------------------
2nd drain DATA:
time = x hrs/job ; rate = 1/x job/hr.
----------------------
EQUATION:
rate + rate = together rate
1/8.5 + 1/x = 1/5
5x + 8.5*5 = 8.5x
3.5x = 42.5
x = 12.14.. hrs
---------------------
Cheers,
Stan H.